OpenStudy (anonymous):

need help with the problem of sqrt 3(4-9)sqrt of 3

OpenStudy (anonymous):

pls draw that out..

OpenStudy (anonymous):

\[\sqrt{3(4-9\sqrt{3}}\]

OpenStudy (anonymous):

\[\sqrt{3(4-9\sqrt{3)}}\] this is the correct one

OpenStudy (anonymous):



OpenStudy (anonymous):

all i can see in this one is distribution. What math level is this?

OpenStudy (anonymous):

someone else gave me this answer and the checkpoint says incorrect

OpenStudy (anonymous):

It says the first step is to use the distributive law and distribute \[\sqrt{3}\]

OpenStudy (anonymous):